# The GenAI Generation: Student Views of Awareness, Preparedness, and Concern

### v1

@@ Abstract
Generative AI (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation: a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines our students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Evaluation of more than 250 responses with more than 40% providing detailed qualitative feedback reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts, with accompanying recommendations to guide educational institutions in navigating a future driven by GenAI.
@@ 1 Introduction
The rapid advancement of artificial intelligence (AI), particularly generative AI (GenAI), is reshaping the educational landscape. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI becomes more embedded in academic and professional contexts, understanding how students perceive its influence on future readiness is essential.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, this research provides actionable insights into how the GenAI Generation, students most affected by GenAI’s rapid integration, understand its role in shaping their academic and career paths.
Defining the GenAI Generation
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [14, 6, 19, 5]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [5, 6]. This study introduces the concept of the GenAI Generation to describe individuals most directly impacted by the Generative AI Revolution. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
@@
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a rapidly changing environment.
@@
This study contributes to the ongoing dialogue about GenAI’s role in shaping students’ academic and professional paths. Centered on awareness, preparedness, and concern, the survey offers a structured view of the GenAI Generation, a group navigating rapid technological change during key years of education. Focusing on higher education, the analysis shows how students weigh GenAI’s potential to support creativity and critical thinking against concerns about overreliance, ethical risks, and reduced engagement. These insights emphasize the need for thoughtful strategies that support ethical and effective use of GenAI in education.
@@ 2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. Designed for completion in under five minutes [16], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [28]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
@@
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [17], this was an intentional design choice to encourage engagement and reduce bias [18]. The survey was administered across two cohorts, with full question text available in the supplemental materials.
@@
The Test 2023 survey was distributed to Computer Science Master’s students via OMSCS (12,000 members) and MSCS (14,000 members) listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1. The Fall 2023 survey was conducted with support from the Office of Academic Effectiveness, which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Methods of Analysis
Survey responses were analyzed using correlation, clustering, and free-response evaluation to capture both quantitative and qualitative insights. These methods supported a systematic exploration of trends and perceptions across diverse student groups.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Degrees were categorized into three groups: Computer Science, Engineering, and Other, based on thematic alignment with disciplinary focus and methodologies. Multidisciplinary fields, such as Robotics and Computational Science and Engineering, were classified under Engineering and Computer Science, respectively, to maintain consistency. This classification enabled cross-disciplinary comparisons and analysis of AI’s impact on skills by degree category and education level.
@@
For the Fall 2023 cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the Test 2023 cohort and 115 for the Fall 2023 cohort
Participation rates varied across groups, with Test 2023 at an estimated 0.5%, and Fall 2023 at 2.5%. These variations may reflect the differences in email dissemination frequency and engagement levels.
Correlation Analysis: To identify relationships between variables, a correlation matrix was computed using the Pearson correlation coefficient, measuring the linear relationship between two variables. The absolute value of 0.5 was the threshold set to identify pairs of features with high collinearity, ensuring that variables with substantial overlap were flagged for further consideration.
@@
Optional Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
@@ 3 Results
The Fall Undergraduates group is divided into three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other, totaling 115 respondents. In contrast, the Test Computer Science Graduates group consists solely of 102 respondents in Computer Science.
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, with CompSci participants slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among Test participants at 100%, followed by CompSci and Engineer respondents at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with Engineers and Test respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Test respondents (93%) and lowest among Engineers (78%).
@@
Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Fall Undergraduate Computer Science students have the highest percentage at 82%, followed by Test Computer Science Graduates at 77%. Fall Other Majors report 56%, while Fall Engineering students have the lowest percentage at 38%.
Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur. Test Computer Science Graduates had the highest expectation of changes occurring the soonest, within 3 years. Fall Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer. The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
@@
Concern: The majority view GenAI’s career impact positively, with Test respondents highest at 81% and Engineering, CompSci, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with optimism and excitement about positive career impact and excitement, concern about GenAI’s impact was also prevalent, with Other disciplines 74% and Engineering 75%, expressing the most caution.
Correlations: Correlations with an absolute value of 0.5 or higher were examined to highlight meaningful patterns in how students view GenAI’s impact. Among Fall undergraduates, excitement about GenAI benefits is moderately correlated (0.51) with perceiving a positive career impact, and satisfaction with GenAI discussions is similarly correlated (0.51) with confidence in professors’ engagement. In computer science, students who believe changes will occur within ten years tend to also view GenAI’s applications as limited, with a correlation of 0.50. Students who feel prepared are more likely to expect positive career outcomes, with a stronger correlation of 0.60. For engineering students, satisfaction with discussions correlates even more strongly with confidence in faculty preparedness, at 0.64. In other disciplines, excitement about GenAI benefits is negatively correlated at -0.57 with the belief that its applications are limited, and positively correlated at 0.76 with a belief in GenAI’s positive career impact. Additionally, satisfaction with discussions is moderately negatively correlated with expecting imminent changes at -0.53 and with anticipating rapid shifts in required skills at -0.55.
Figure 5: Topics raised in free response
Free Response: Optional free response rates were relatively high, with 40–50% of undergraduates and test graduate students responding to the institutional recommendation prompt, and 23–25% responding to the additional thoughts prompt. Due to limited responses by discipline, the Fall undergraduate cohort was analyzed as a whole.
The most common recommendation—mentioned by over 50% of both cohorts—was teaching students how to use GenAI tools. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses.
Emotional tone varied between cohorts. Over half of undergraduates expressed negative emotions, 29% more than the test graduate group, while test graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes. Test graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
@@ 4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers.
@@
While overall awareness of GenAI is high, this awareness does not translate uniformly into understanding or confidence. Non-STEM students exhibited the strongest correlation between excitement about GenAI and its perceived career impact, suggesting an optimistic yet potentially less critically informed perspective. In contrast, engineering students had the highest skepticism about GenAI’s practical applications, while computer science students were most likely to view GenAI as a short-lived trend. These viewpoints may reflect the varied depth of engagement with AI technologies in their respective fields, with technical students balancing enthusiasm with caution about GenAI’s limitations and ethical implications.
These nuances are echoed in free responses, where students shared diverse sentiments:
@@
For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
@@
While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead. Unlike future students who may benefit from GenAI-integrated curricula, the current generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills. Among computer science students, there is a clear pattern: those who feel prepared are most likely to expect a positive career impact.
In contrast, only 56% of ”Other” students and 38% of engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
@@
Correlations support these patterns. In engineering, classroom discussions on GenAI are linked to higher confidence in professors, suggesting that open engagement builds trust and a sense of readiness. In computer science, skepticism often aligns with delayed expectations, while students who feel prepared tend to view GenAI as a career asset. In the ”Other” cohort, satisfaction with discussions correlates with the belief that GenAI will impact job skills, suggesting confidence in the status quo or comfort with limited exposure.
The free response data adds depth to these findings. Despite overall confidence in their curriculum, over half of students recommended more courses on GenAI and its practical use. Many also called for stronger emphasis on ”fundamentals,” ”critical thinking,” and ethical considerations:
@@
The findings reveal a misalignment between perceived career impact and readiness. As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. Addressing this gap will require expanding GenAI education across fields, embedding GenAI literacy into non-STEM programs, and equipping faculty to lead informed, inclusive discussions. These efforts are essential to preparing a more equitable and capable generation for a GenAI-driven future.
Concern: learning at what cost?
Student concerns about GenAI range from ethics and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
@@
“im excited for the changes and technological advances it will bring.“ — Graduate Computer Science
@@
Free responses reflect strong concern, especially in the “additional thoughts” section. Over 50% conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
@@
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Institutions must actively support both students and faculty as they navigate this rapidly evolving landscape.

### v2

@@ Abstract
Generative Artificial Intelligence (GenAI) is revolutionizing education and workforce development, profoundly shaping how students learn, engage, and prepare for their future. Outpacing the development of uniform policies and structures, GenAI has heralded a unique era and given rise to the GenAI Generation. We define the GenAI Generation as a cohort of students whose education has been increasingly shaped by the opportunities and challenges GenAI presents during its widespread adoption within society. This study examines students’ perceptions of GenAI through a concise survey with optional open-ended questions, focusing on their awareness, preparedness, and concerns. Notably, readiness appears increasingly tied to exposure to GenAI through one’s coursework. Students with greater curricular exposure to GenAI tend to feel more prepared, while those without it more often express vulnerability and uncertainty, highlighting a new and growing divide in readiness that goes beyond traditional disciplinary boundaries. Evaluation of more than 250 responses, with over 40% providing detailed qualitative feedback, reveals a core dual sentiment: while most students express enthusiasm for GenAI, an even greater proportion voice a spectrum of concerns about ethics, job displacement, and the adequacy of educational structures given the highly transformative technology. These findings offer critical insights into how students view the potential and pitfalls of GenAI for future career impacts. The challenge ahead involves implementing associated recommendations for educational institutions, moving beyond the baseline of access toward more informed guidance on the use of these tools, while preserving critical thinking, ethical reasoning, and adaptive learning.
Keywords: Generative Artificial Intelligence; GenAI Generation; Student Perceptions; Higher Education; Career Readiness; Cognitive Development; Adaptability; Problem-Solving; Critical Thinking; Ethics in AI
@@ 1 Introduction
The significant advancement of Artificial Intelligence (AI), particularly Generative AI (GenAI), is reshaping the educational landscape. GenAI is a type of AI trained on large datasets to generate new content such as text, images, audio, or video in response to user prompts, setting it apart from predictive models by its ability to create [22, 28, 20, 21]. As a tool capable of autonomously producing human-like content, GenAI offers both opportunities and uncertainties for education. It can enhance creativity, streamline workflows, and support personalized learning, yet its broader impact remains unclear. As GenAI continues to evolve exponentially and integrate within academic contexts, there is a pressing need to understand how it is shaping the developmental experiences of students in their formative academic years. Unlike previous waves of technology’s impact on education, GenAI’s pace and scale introduce a layer of uncertainty around its long-term effects on cognition, critical thinking, and engagement. By honing in on student perspectives, this research provides early insight into how GenAI is influencing not only perceived career readiness but also the cognitive processes underpinning learning and decision-making.
This study presents the development, dissemination, and evaluation of a survey examining student perspectives on GenAI through three core lenses: awareness, preparedness, and concern. By analyzing responses, it provides actionable insights into how the GenAI Generation—students most directly impacted by the rise of GenAI—understand its role in shaping both their academic experiences and future trajectories.
The GenAI Generation
The GenAI Revolution represents a paradigm shift in technology adoption, enhancing efficiency, creativity, and performance while introducing challenges such as ethical concerns and the need for responsible stewardship [11, 5, 16, 4]. As it relates to education, this revolution not only emphasizes a transformation in learning but also highlights the challenges students and teachers face during a phase of unclear uniform guidance [4, 5]. This study introduces the GenAI Generation. The GenAI Generation is defined as the cohort of individuals actively undergoing developmental and educational growth/attainment during the rise of Generative AI, starting with the release of ChatGPT in November 2022. The GenAI Generation encompasses the experiences of students from high school through higher education, during the initial, transformative years of the GenAI Revolution, when the integration of AI tools is still experimental, unregulated, and highly influential. Unlike earlier generations who did not have these tools and later generations who may inherit more structured systems, this group faces unique challenges and opportunities surrounding uncertainty.
@@
  • High Adaptability: Balancing both the traditional and new competencies required to effectively interact and engage with GenAI in a highly dynamic environment.
@@
This study offers timely insights into how students in their formative educational years are experiencing the rapid rise of Generative AI. It introduces the term “GenAI Generation” to describe students navigating both academic development and technological transformation without consistent institutional guidance or established norms.
Drawing from a combination of structured survey data and rich qualitative responses, the study reveals meaningful patterns in students’ awareness, preparedness, and concern regarding GenAI. Across majors, students expressed a desire for more intentional support, not just access to GenAI tools, but structured opportunities to learn how to use them responsibly. While formal curriculum integration remains inconsistent, students who encounter GenAI through coursework or self-directed exploration are more likely to articulate confidence and optimism about its role in their future.
From enthusiasm to anxiety, the findings illuminate the psychosocial and pedagogical dimensions of GenAI’s presence in higher education. Students are not only engaging with a new technology, they are forming perceptions of their own readiness, trust in institutions, and ethical responsibility in real time. These perspectives carry clear implications for curriculum design, faculty preparation, and equitable access to digital literacy.
This work contributes to an emerging understanding of what students need to thrive in a GenAI-influenced world: not only technical fluency, but critical engagement, ethical grounding, and institutional clarity. The study underscores the importance of proactive educational responses that extend beyond STEM fields to support students across all disciplines in this pivotal moment.
@@ 2 Methods
The survey targeted students within the GenAI Generation, focusing on GenAI’s perceived impact on their education and careers. Awareness, preparedness, and concern guided the development of eight main questions, three demographic items, and two open-ended prompts. This study received Institutional Review Board (IRB) approval from the [Institution Name Redacted for Review], and all participants provided informed consent prior to beginning the survey. Designed for completion in under five minutes [13], the survey began with an IRB-required disclosure and a screening question, followed by items progressing from general to sensitive topics [24]. Questions were refined through lab feedback, aligned with key concepts, and included internal validation to ensure consistency and quality.
@@
A four-point, forced-choice Likert-type scale was used to assess attitudes, significance, intensity, and satisfaction. While omitting a neutral option is debated [14], this was an intentional design choice to encourage engagement and reduce bias [15]. The survey was administered across two [Institution Name Redacted for Review] student cohorts, with full question text available in the supplemental materials.
The first cohort, Fall 2023 Computer Science Graduate students, survey was distributed to Online Computer Science Master’s (OMSCS) students, approximately 12,000 members [CITE REDACTED FOR REVIEW], and campus Computer Science Master’s (MSCS), approximately 14,500 members [CITE REDACTED FOR REVIEW], via listservs. Invitations were sent on September 28 and 29, 2023, with responses collected through November 1.
@@
The second cohort, Fall 2023 Undergraduate students, survey was conducted with support from the [Institution Name Redacted for Review] Office of Academic Effectiveness (OAE), which used stratified random sampling to select 5,655 undergraduates representative of the broader student population. Sampling quotas ensured balance by gender, race, citizenship, and class standing, with a 2% margin of error. The demographic breakdown of the sample can be found in the supplemental materials. The survey was open from November 2–23, 2023, with email reminders sent on November 2, 15, and 22.
Analysis Approach
Descriptive statistics and frequency distributions were used to compare response patterns across student subgroups, while qualitative evaluation of open-ended responses enabled a systematic exploration of emerging themes and perceptions.
Data Processing: To ensure data integrity, only completed responses containing degree type information were included in the analysis. Where the first cohort of graduate students comprised soley of computer science majors, the undergraduate cohort had many degrees. Rather than broadly distinguishing between STEM and non-STEM, degrees were categorized into Computer Science, Engineering, and Other, based on disciplinary focus and methodological orientation. Table 1 provides a Major-Category mapping for reference. This grouping structure reflects both practical sample sizes (Other, 28 Computer Science, and 60 Engineering) and a meaningful distinction in how students may engage with GenAI in academic contexts. This also enabled a comparison between computer science graduate and undergraduate perspectives.
Grouping a diverse range of degrees—such as Biology, Business Administration, Literature, and Public Policy—under the single category “Other” introduces a potential limitation, as it may obscure important disciplinary differences in students’ exposure to or perspectives on GenAI. While specific exposure to AI tools and curricula was not measured directly, these categories serve as a reasonable proxy for varying degrees of disciplinary integration with GenAI-related content, allowing exploratory comparisons across educational domains.
Table 1: Undergraduate Sample and Response Rates by Majors and Category
Major* Category Sample Freq. Response Freq. Sample % Response %
Applied Physics Other 12 2 0.2 1.7
Biology Other 162 3 2.9 2.6
Business Administration Other 440 3 7.8 2.6
Computational Media Other 98 1 1.7 0.9
Earth & Atmospheric Sciences Other 13 1 0.2 0.9
Econ & Int’l Affairs Other 15 1 0.3 0.9
Economics Other 49 2 0.9 1.7
History, Technology, & Society Other 18 2 0.3 1.7
Industrial Design Other 88 2 1.6 1.7
Int’l Affairs & Mod Lang Other 18 2 0.3 1.7
Lit., Media, & Communication Other 59 2 1.0 1.7
Mathematics Other 77 3 1.4 2.6
Neuroscience Other 146 3 2.6 2.6
Physics Other 68 5 1.2 4.3
Psychology Other 66 5 1.2 4.3
Public Policy Other 37 1 0.7 0.9
Aerospace Engineering Engineer 349 8 6.2 7.0
Biomedical Engineering Engineer 399 6 7.1 5.2
Chemical and Biomolecular Eng Engineer 187 5 3.3 4.3
Computer Engineering Engineer 187 2 3.3 1.7
Electrical Engineering Engineer 187 3 3.3 2.6
Environmental Engineering Engineer 58 3 1.0 2.6
Industrial Engineering Engineer 344 5 6.1 4.3
Materials Science & Engr Engineer 67 3 1.2 2.6
Mechanical Engineering Engineer 582 12 10.3 10.4
Nuclear & Radiological Engr Engineer 32 2 0.6 1.7
Computer Science Computer Science 1419 28 25.1 24.3
Applied Lang/Intercultural St No Response 4 0 0.1 0.0
Architecture No Response 76 0 1.3 0.0
Atmospheric & Oceanic Sciences No Response 3 0 0.1 0.0
Biochemistry No Response 88 0 1.6 0.0
Building Construction No Response 13 0 0.2 0.0
Chemistry No Response 40 0 0.7 0.0
Environmental Science No Response 10 0 0.2 0.0
Global Econ/Mod Lang No Response 5 0 0.1 0.0
International Affairs No Response 22 0 0.4 0.0
Music Technology No Response 23 0 0.4 0.0
Solid Earth/Planetary Sciences No Response 2 0 0.0 0.0
Special/Non-Degree No Response 64 0 1.1 0.0
Civil Engineering No Response 128 0 2.3 0.0
*For reference; sample was not balanced on this attribute. Sample breakdowns by additional attributes can be found within the supplemental material.
@@
For the undergraduate cohort, two non-undergraduate entries, likely from dual-enrolled students, were excluded to maintain a focus on undergraduate perspectives. The remaining responses totals 102 for the graduate cohort and 115 for the undergraduate cohort. Table 2 shows this process
Table 2: Table showing data cleaning process
Survey Population Total Started Total Incomplete Total Dropped Reason for Drop Remaining Completed Total
Computer Science Graduates 2023 26,000** 122 19 1 Response was “prefer not to answer” for degree 102
Undergraduates 2023 5,655 143 26 2 Did not have “undergraduate” for degree type 115
** Approximate total population estimate for Computer Science Graduate 2023 cohort.
@@
Qualitative Free-Response Analysis: An inductive coding approach was used to analyze free-response data. Preprocessing included tokenization, lowercasing, and the removal of punctuation and stop words. N-grams (unigrams to quadgrams) were extracted, with low-frequency patterns excluded. Grammatical parsing identified noun and verb phrases, which, alongside n-grams, helped surface recurring themes and patterns.
@@ 3 Results
The undergraduate cohort included 115 respondents across three categories: 28 in Computer Science, 60 in Engineering, and 27 in Other disciplines. The graduate cohort response count was 102 and comprised solely of computer science majors. The full breakdown between sample to response counts can be seen in Table 1.
Figure 2: Response counts by degree and level
Awareness: A small percentage of respondents in all groups believe GenAI is a short-lived trend, seen in Figure 3, with undergraduate computer science students slightly more likely to agree at 21%. Most respondents view generative AI as a significant advancement, with nearly unanimous agreement among the graduate cohort at 100%, followed by undergraduate computer science and engineers at 96% and 93%, respectively. When asked about the limitations of generative AI’s potential applications, agreement levels vary, with engineers and graduate respondents showing slightly higher agreement at 34% and 37%. Generative AI experience is highest among Graduate respondents (93%) and lowest among Engineers (78%).
@@
Preparedness: For the percentage of respondents that believe job skills will change due to GenAI, Figure 4 shows undergraduate computer science students have the highest percentage at 82%, followed by computer science graduates at 77%. Other majors report 56%, while engineering students have the lowest percentage at 38%.
Figure 4: Responses on if job skills will change due to GenAI.
Of those that answered “Yes” to job skills changing, there was a follow up question of how soon they believe the changes will occur. Figure 5 shows computer science graduates had the highest expectation of changes occurring the soonest. Undergraduate Computer Science and Other Majors have a more balanced distribution, with nearly half predicting changes within 3 years and the rest expecting shifts within 5 years or longer. The Fall Undergraduate Engineering students mostly expect changes within 5 or 10 years and just 17% expected changes to occur within 3 years.
@@
Concern: The majority view GenAI’s career impact positively, seen in Figure 7, with computer science graduates highest at 81% and undergraduate Engineers, Computer Science, and Other lower with 70%, 64%, and 63% respectively. Excitement about GenAI benefits was high across all groups, peaking among Engineering at 77% and Test at 80%. Other respondents notably lower at 67%. Even with excitement and optimism about positive career impacts, concern was even greater across all undergraduate degree categories. While less, the computer science graduate cohort still showed concern with 65% concerned with GenAI’s.
Figure 7: Responses on concern
Free Response: In addition to the structured survey responses, a notable portion of participants contributed to the optional open-ended questions. These qualitative insights deepen understanding of the perspectives behind the patterns and are summarized below. 40–50% of undergraduates and graduate students responded to the institutional recommendation prompt, and 23–25% provided input on the additional thoughts prompt. Participation rates for the free response can be seen in Table 3, showing non-Computer Science students having the highest participation. The Fall undergraduate free responses were analyzed as a whole.
Table 3: Optional Free Response Participation by Question & Degree Category
Category Institutional Recommendation Total Additional Thoughts Total Survey Response Total
Undergraduate Other 17 11 38
Undergraduate Engineer 24 11 49
Undergraduate Computer Science 6 5 29
Graduate Computer Science 52 27 102
Figure 8: Topics raised in free response for institutional recommendations
The most common recommendation, mentioned by over 50% for both cohorts, was teaching students how to use GenAI tools (Figure 8. About 33% of respondents also suggested integrating GenAI into existing curricula or adding new courses. Ethics and career implications of GenAI appeared more frequently in undergraduate responses, with ethics mentioned 15% more and career impact 5% more than in graduate responses. There were fewer who responded to the optional ”additional thoughts” question, those who did showed clear emotional tone(Figure 9). Over half of undergraduates expressed negative emotions, 29% more than the graduate cohort, while graduates were 20% more likely to express positivity. Additional thoughts covered a broader range of themes(Figure 10). The computer science graduates frequently emphasized critical thinking, whereas undergraduates more often expressed uncertainty.
Figure 9: Percentage of free responses with either positive or negative emotions conveyed
Figure 10: Topics raised in free response additional thoughts
@@ 4 Discussion
The results have presented unique and shared perspectives across undergraduate programs, revealing areas of improvement. Thanks to the participants in this study and their contributions to the free response, there is a story that has begun to emerge surrounding the needs of the GenAI Generation as they learn and grow into their future careers. Although surveys were distributed uniformly to ensure equal access across cohorts, the potential for selection bias and low response frequencies remains a limitation. Response counts were modest compared to sample size, and some subgroups—particularly within non-CS or less tech-aligned majors—were underrepresented. As a technology-focused institution, [Institution Name Redacted for Review]’s student population may also have greater exposure to emerging technologies, including GenAI, than students at other institutions. Even with these constraints, the findings offer important early insight into how GenAI is shaping student experiences, perceptions, and expectations during a critical period of educational and technological transformation.
@@
While overall awareness of GenAI was high across all groups, levels of understanding and confidence varied. Computer science undergraduates were the most likely to consider GenAI a short-lived trend (21%), whereas engineering and “Other” majors reported lower agreement (10–11%). Despite this skepticism, nearly all respondents across disciplines viewed GenAI as a meaningful advancement, with agreement ranging from 89% to 100%. Experience using GenAI tools also varied, with graduate students reporting the highest exposure (93%) and engineering undergraduates the lowest (78%). Responses regarding the limitations of GenAI’s potential were more common among engineering and graduate students. For the graduate cohort comprising of computer science, this could suggest increased exposure may come with greater critical awareness of its boundaries. For the undergraduate engineering students, skepticism may come from the belief that GenAI’s application is limited specifically for their degree. These nuances are echoed in free responses, where students shared diverse sentiments:
@@
These emotional responses and findings reveal a misalignment between perceived career impact and readiness and signal a broader psychosocial landscape in flux where students must construct academic identity and career expectations while navigating an unregulated and evolving technology. For a transformative tool like GenAI, students are actively seeking guidance to understand its potential, limits, and ethical implications.
@@
While students recognize GenAI’s growing influence on their careers, many remain unsure whether their education is preparing them for the changes ahead. Unlike future students who may benefit from GenAI-integrated curricula, the GenAI Generation often relies on self-directed strategies to prepare for a GenAI-driven workforce.
This preparedness gap is reflected in how different groups perceive GenAI’s impact on job skills. There is a clear pattern that begins to emerge: those who feel prepared in their degree are most likely to expect a positive career impact.
Where computer science 77% of graduates and 82% undergraduates anticipate job skill changes, only 56% and 38% of undergaduate other and engineering students anticipate such changes, with many engineering students expecting a slower pace—highlighting a lack of urgency in their field. These differences suggest the need for tailored strategies that align with each group’s career context.
@@
The free response data adds depth to these findings. Despite overall confidence in their curriculum, in responding to the first institutional recommendations response question, over half of students wanted to be taught how to use GenAI. Many also called for stronger emphasis on fundamentals, critical thinking, and ethical considerations:
@@
As GenAI begins to enter technical curricula, students in other disciplines risk being left behind. This gap underscores a fundamental pedagogical challenge: while students recognize GenAI’s relevance, they often do not see it meaningfully integrated into their learning environments, leaving students to independently navigate tools that carry both potential and risk. This lack of structure points to an urgent need for faculty guidance and training, intentional curriculum design, and interdisciplinary GenAI literacy across academic programs. These efforts are essential to prepare a more equitable and capable generation for a GenAI-driven future.
Concern: use at what cost?
Student concerns about GenAI range from ethical implications and job security to institutional and faculty preparedness. While excitement is widespread, it is tempered by deeper anxieties, reflecting the complexity of the GenAI Generation’s experience. The Test 2023 Computer Science Graduate cohort reported the lowest concern (65%) and the highest levels of excitement (80%) and belief in GenAI’s positive career impact (81%). This likely reflects the integration of GenAI into their curriculum and their confidence using such tools.
@@
“I’m excited for the changes and technological advances it will bring.“ — Graduate Computer Science
@@
Free responses reflect strong concern, especially in the the open ended qustion asking for additional thoughts. Over 50% of undergraduates conveyed negative emotions, with recurring themes of ethics, job replacement, and uncertainty—often expressed in long, passionate comments.
@@
A compelling insight lies in who chose to share their perspectives in the optional free-response section. Among undergraduates, participation was highest among engineering and “other” majors, with response rates of 50% and 44%, respectively-compared to just 20% among computer science undergraduates. This could suggest that students who may not encounter GenAI in their formal coursework still feel strongly about its impact and relevance.
Student responses reveal strong concerns about the ethical, economic, and societal impacts of GenAI, alongside a desire for greater guidance and critical engagement. Students are not only grappling with how to use GenAI but also with what kind of future it is shaping.